# Los Angeles, CA, USA AtomGS

### v2

@@ 4.1 Atomized Proliferation
Atom Gaussian Primitive: Our process begins by analyzing the input SfM points to establish the Atom scale $\mathcal{S}_{a}=P_{1}(\bm{d})$, calculated from the first percentile of ordered distances. This scale distinguishes between Gaussians capturing fine details (Atom Gaussians) and those covering broader background elements (traditional Gaussians). Atom Gaussians are distinct in being isotropic spheroids with a uniform size ($\bm{s}_{1}=\bm{s}_{2}=...=\bm{s}_{i}=(\mathcal{S}_{a})^{3}$), in contrast to traditional Gaussians, which are anisotropic ellipsoids of varying sizes. The constant size of Atom Gaussians imposes the priority of densifying them to accurately fill gaps in the geometry, rather than optimizing large Gaussians that merely cover these voids. The uniform size of Atom Gaussians ensures a closer alignment with the actual 3D geometry of the scene, leading to more accurate surface representations.
Atomization: This step checks the condition $||\bm{S}||_{\min}\leq\mathcal{S}_{a}$ regularly. If met, the Gaussian is designated as an Atom Gaussian with a size set to $\bm{s}=(\mathcal{S}_{a})^{3}$. Once categorized as Atom Gaussians, their scales are fixed and no longer optimized through backpropagation but through a geometric progression: $\mathcal{S}_{an}=\mathcal{S}_{an}\cdot r^{n-1}$, where $r=\sqrt[t_{a}]{p}$ and $t_{a}$ represents the total number of atomization iterations, and $p$ is the final proportion to the initial $\mathcal{S}_{a}$. This ensures that the atom scale decreases over iterations, progressively enhancing the representation of fine details.
@@ 4.2 Geometry-Guided Optimization
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Normal (b) Curvature (c) Edge (d) Normal Loss
Figure 3: Illustration of Edge-Aware Normal Loss. The normal map is rendered from the 3DGS 30k result. Based on that, we compute the normal loss map, showing the areas that our proposed loss could optimize.
@@
$\mathcal{L}=(1-\lambda_{SSIM})\mathcal{L}_{1}+\lambda_{SSIM}\mathcal{L}_{MS-SSIM}+\lambda_{Normal}\mathcal{L}_{Normal},$ (3)
where $\lambda_{SSIM}$ and $\lambda_{Normal}$ are hyperparameters that determine the respective contributions of $\mathcal{L}_{MS-SSIM}$ and $\mathcal{L}_{Normal}$ to the overall loss function.
@@ 5 Experiments
3D Geometry precision: Table 2 presents a quantitative comparison between our approach and other methods aimed at improving 3D reconstruction accuracy. Our approach not only surpasses other explicit methods in geometric precision but also competes favorably with SDF-based implicit methods. Note that compared with other implicit methods, our approach also benefits from faster training speeds, thereby increasing its practical applicability in real-world scenarios. Figure 5 presents a qualitative comparison of mesh reconstruction using 3DGS, SuGaR, and our method on DTU scene 24. SuGaR tends to generate flat disks due to its regularization approach; however, these disks are not always perfectly aligned with the underlying geometry. NeuS employs a signed distance function, resulting in smoother surfaces, but could also lose high-frequency details due to strong smoothness priors. In contrast, our method maintains a balance between smoothness and detail preservation.
Ablation Study: We conducted three ablation studies on the Tanks & Temples dataset [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] to evaluate the impact of different configurations on our model’s performance. The first configuration involved the removal of Atomized Proliferation from our model. Following that, we kept Atomized Proliferation but discarded either of the loss functions during training. The results from these configurations and the full model are shown in Table 3. The results indicate that Atomized Proliferation is crucial to our model’s performance. Excluding Atomized Proliferation resulted in the most significant performance decline. This suggests that simply incorporating geometry regularization in the loss functions tend to merely approximate the scene with elongated Gaussian ellipsoids instead of aligning accurately with the underlying geometry. Additionally, when each of the two advanced loss components was individually removed, there was a slight decrease in performance. This demonstrates the importance of each component in achieving optimal results.
@@
Refer to caption Refer to caption Refer to caption
(a) Ours (b) NeuS (c) SuGaR
Figure 5: Mesh Comparison on the DTU Dataset.
Tanks&Temples
$\text{PSNR}^{\uparrow}$ $\text{SSIM}^{\uparrow}$ $\text{LPIPS}^{\downarrow}$
No $atomization$ 23.25 0.815 0.228
No $L_{normal}$ 23.58 0.840 0.181
No $L_{ms-ssim}$ 23.58 0.837 0.185
Full model 23.70 0.849 0.166
Table 3: Ablation Study.
@@ 6 Conclusion
Nonetheless, AtomGS has its own limitations. Similar to the previous methods, our method may not produce accurate geometry for highly specular or semi-transparent material. While our method in general requires fewer GS primitives than the original 3DGS method to achieve improved visual quality, our proliferation strategy could sometimes produce more GS primitives to represent all details for highly complex environments. In the future, we aim to develop an improved pruning strategy to achieve a more compact result.
@@ Appendix A Atomized Proliferation Algorithm
The Atomized Proliferation algorithm is summarized in Algorithm 1. It starts with setting parameters for Clone Threshold ($\tau_{c}$), Split Threshold ($\tau_{s}$), Prune Threshold ($\epsilon$), Atom Scale ($\mathcal{S}_{a}$), and defining duration limits for atomized proliferation ($t_{a}$) and warm-up phase ($t_{w}$). The algorithm iteratively processes each Gaussian property ($\bm{\mu},\bm{\Sigma},\bm{c},\alpha$) from the Gaussian set ($\bm{M},\bm{S},\bm{C},\bm{A}$). A Gaussian is pruned if its $\alpha$ falls below the threshold $\epsilon$ or its covariance ($\bm{\Sigma}$) is excessively large. If the gradient of the loss ($\nabla_{p}\mathcal{L}$) exceeds $\tau_{c}$, the Gaussian is cloned to potentially bridge geometry gaps. Additionally, the Gaussian is split when $\nabla_{p}\mathcal{L}$ meets a dynamically adjusted threshold based on the warm-up progress and if the norm of $\bm{\Sigma}$ exceeds the Atom Scale ($\mathcal{S}_{a}$). Atomization takes place when the minimum norm of $\bm{S}$ is less than or equal to Atom Scale $\mathcal{S}_{a}$ and within the proliferation timeframe ($t_{a}$), ensuring detail refinement before the proliferation endpoint.
@@
0:  Clone Threshold $\tau_{c}$, Split Threshold $\tau_{s}$, Prune Threshold $\epsilon$,     Atom Scale $\mathcal{S}_{a}$, Atomized Proliferation until $t_{a}$, Warm-Up until $t_{w}$
@@
8:     if $\nabla_{p}\mathcal{L}\geq\min\left(\frac{i}{t_{w}}\tau_{s},\tau_{s}\right)$ and $||\bm{S}||_{\max}>\mathcal{S}_{a}$ then
@@ Appendix C Implementation Details
Hyper Parameter Settings: Following the 3DGS, we set the Clone Threshold at $\tau_{c}=0.002$, Split Threshold at $\tau_{s}=0.002$, and Prune Threshold at $\epsilon=0.005$. For the Atom-related settings, we set Atom Scale at the first percentile of distances from the input SfM points $\mathcal{S}_{a}=P_{1}(\bm{d})$, Atomized Proliferation until iteration at $t_{a}=7000$, and Warm-Up until iteration at $t_{w}=7000$. For optimization, the weights for SSIM and normal calculations are both set at $\lambda_{SSIM}=\lambda_{Normal}=0.1$. When working with object-centered datasets that lack extensive backgrounds, we advise setting the scale learning rate $\eta_{s}=0$ to maximize geometric accuracy.
@@ Appendix E Additional Results
Figure 8: Radiance Field Comparison on the Mip-NeRF360 Dataset.
@@
DTU 24
Refer to caption Refer to caption Refer to caption
@@
Hotdog
Refer to caption Refer to caption Refer to caption
@@
Figure 9: Mesh Comparison on DTU Dataset and NeRF Synthetic Dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]

### v3

@@ 4.1 Atomized Proliferation
Atom Gaussian Primitive: Our process begins by analyzing the input SfM points to establish the Atom scale $\mathcal{S}=P_{i}(\bm{d})$, calculated from the $i$th percentile of ordered distances. This scale distinguishes between Gaussians capturing fine details (Atom Gaussians) and those covering broader background elements (traditional Gaussians). Atom Gaussians are distinct in being isotropic spheroids with a uniform size ($\bm{s}_{1}=\bm{s}_{2}=...=\bm{s}_{n}=(\mathcal{S})^{3}$), in contrast to traditional Gaussians, which are anisotropic ellipsoids of varying sizes. The uniform size of Atom Gaussians imposes the priority of densifying them to accurately fill gaps in the geometry and ensures a closer alignment with the actual 3D geometry of the scene, rather than optimizing elongated Gaussians that approximately cover these voids.
Atomization: This step checks the condition $||\bm{S}||_{\min}\leq\mathcal{S}$ regularly. If met, the Gaussian is designated as an Atom Gaussian with a size set to $\bm{s}=(\mathcal{S})^{3}$. Once categorized as Atom Gaussians, their scales are fixed and no longer optimized through backpropagation but through a geometric progression: $\mathcal{S}_{n}=\mathcal{S}_{n-1}\cdot r^{n-1}$, where $r=\sqrt[t_{a}]{p}$ and $t_{a}$ represents the total number of atomization iterations, and $p$ is the final proportion to the initial $\mathcal{S}$. This ensures that the atom scale decreases over iterations, progressively enhancing the representation of fine details.
@@ 4.2 Geometry-Guided Optimization
Refer to caption Refer to caption
(a) Normal (b) Curvature
Refer to caption Refer to caption
(c) Edge (d) Edge-Aware Normal Loss
Figure 3: Illustration of the Edge-Aware Normal Loss. The image maps are rendered from the 3DGS 30k result without applying our proposed Edge-Aware Normal Loss. Loss map (d) highlights the areas where it can improve the geometry over the 3DGS result.
@@
$\mathcal{L}=(1-\lambda_{ms-ssim})\mathcal{L}_{1}+\lambda_{ms-ssim}\mathcal{L}_{ms-ssim}+\lambda_{normal}\mathcal{L}_{normal},$ (3)
where $\lambda_{ms-ssim}$ and $\lambda_{normal}$ are hyperparameters that determine the respective contributions of $\mathcal{L}_{ms-ssim}$ and $\mathcal{L}_{normal}$ to the overall loss function.
@@ 5 Experiments
DTU 24
Refer to caption Refer to caption Refer to caption
Hotdog
Refer to caption Refer to caption Refer to caption
(a) Ours (b) NeuS (c) SuGaR
Figure 5: Mesh Comparison on the DTU and NeRF Synthetic Datasets [Aanæs et al.(2016)Aanæs, Jensen, Vogiatzis, Tola, and Dahl, Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]
@@
3D Geometry precision: Table 2 presents a quantitative comparison between our approach and other methods aimed at improving 3D reconstruction accuracy. Our approach not only surpasses other explicit methods in geometric precision but also competes favorably with SDF-based implicit methods. Note that compared with other implicit methods, our approach also benefits from faster training speeds, thereby increasing its practical applicability in real-world scenarios. Figure 5 presents a qualitative comparison of mesh reconstruction using 3DGS, SuGaR, and our method on DTU scenes 24, 106, and 122. SuGaR tends to generate flat disks due to its regularization approach; however, these disks are not always perfectly aligned with the underlying geometry. NeuS employs a signed distance function, resulting in smoother surfaces, but could also lose high-frequency details due to strong smoothness priors. In contrast, our method maintains a balance between smoothness and detail preservation.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) No $atomization$ (b) No $\mathcal{L}_{normal}$ (c) No $\mathcal{L}_{ms-ssim}$ Full model
Figure 6: Qualitative Results of Ablation on the DTU scene 24.
Ablation Study: We conducted ablation studies on the Tanks & Temples and DTU datasets [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun, Aanæs et al.(2016)Aanæs, Jensen, Vogiatzis, Tola, and Dahl] to evaluate the impact of different configurations on our model’s performance. The first configuration involved the removal of Atomized Proliferation from our model. Following that, we kept Atomized Proliferation but discarded either of the loss functions during training. The results from these configurations and the full model are shown in Table 3 and Figure 6. The results indicate that Atomized Proliferation is crucial to our model’s performance. Excluding Atomized Proliferation resulted in the most significant performance decline. This suggests that simply incorporating geometry regularization in the loss functions tends to merely approximate the scene with elongated Gaussian ellipsoids instead of aligning accurately with the underlying geometry. Additionally, omitting either $\mathcal{L}_{normal}$ loss or $\mathcal{L}_{ms-ssim}$ loss also leads to performance degradation, but to a lesser extent. The full model, which incorporates all components, consistently achieves the best performance, indicating that each element contributes to the overall effectiveness of the model.
Tanks&Temples DTU
[…]
No $atomization$ 23.25 0.815 0.228 1.23
No $\mathcal{L}_{normal}$ 23.58 0.840 0.181 0.82
No $\mathcal{L}_{ms-ssim}$ 23.58 0.837 0.185 0.78
Full model 23.70 0.849 0.166 0.75
Table 3: Quantitative Results of Ablation on the Tanks&Temples and DTU datasets.
@@ 6 Conclusion
Nonetheless, AtomGS has its own limitations. Similar to the previous methods, our method may not produce accurate geometry for highly specular or semi-transparent material. While our method in general requires fewer GS primitives than the original 3DGS method to achieve improved visual quality, our proliferation strategy could sometimes produce more GS primitives to represent all details for highly complex environments. In the future, we aim to develop an improved pruning or merging strategy to achieve a more compact result.
Acknowledgments
The authors would like to thank our primary sponsors of this research: Mr. Clayton Burford of the Battlespace Content Creation (BCC) team at Simulation and Training Technology Center (STTC). This work is supported by University Affiliated Research Center (UARC) award W911NF-14-D-0005. Statements and opinions expressed and content included do not necessarily reflect the position or the policy of the Government, and no official endorsement should be inferred.
@@ Appendix A Atomized Proliferation Algorithm
The Atomized Proliferation algorithm is summarized in Algorithm 1. It starts with setting parameters for Clone Threshold ($\tau_{c}$), Split Threshold ($\tau_{s}$), Prune Threshold ($\epsilon$), Atom Scale ($\mathcal{S}$), and defining duration limits for atomized proliferation ($t_{a}$) and warm-up phase ($t_{w}$). The algorithm iteratively processes each Gaussian property ($\bm{\mu},\bm{\Sigma},\bm{c},\alpha$) from the Gaussian set ($\bm{M},\bm{S},\bm{C},\bm{A}$). A Gaussian is pruned if its $\alpha$ falls below the threshold $\epsilon$ or its covariance ($\bm{\Sigma}$) is excessively large. If the gradient of the loss ($\nabla_{p}\mathcal{L}$) exceeds $\tau_{c}$, the Gaussian is cloned to potentially bridge geometry gaps. Additionally, the Gaussian is split when $\nabla_{p}\mathcal{L}$ meets a dynamically adjusted threshold based on the warm-up progress and if the norm of $\bm{\Sigma}$ exceeds the Atom Scale ($\mathcal{S}$). Atomization takes place when the minimum norm of $\bm{S}$ is less than or equal to Atom Scale $\mathcal{S}$ and within the proliferation timeframe ($t_{a}$), ensuring detail refinement before the proliferation endpoint.
@@
0:  Clone Threshold $\tau_{c}$, Split Threshold $\tau_{s}$, Prune Threshold $\epsilon$,     Atom Scale $\mathcal{S}$, Atomized Proliferation until $t_{a}$, Warm-Up until $t_{w}$
@@
8:     if $\nabla_{p}\mathcal{L}\geq\min\left(\frac{i}{t_{w}}\tau_{s},\tau_{s}\right)$ and $||\bm{S}||_{\max}>\mathcal{S}$ then
@@ Appendix C Implementation Details
Hyper Parameter Settings: Following the 3DGS, we set the Clone Threshold at $\tau_{c}=0.002$, Split Threshold at $\tau_{s}=0.002$, and Prune Threshold at $\epsilon=0.005$. For the Atom-related settings, we set Atom Scale at the first percentile of distances from the input SfM points $\mathcal{S}=P_{1}(\bm{d})$, Atomized Proliferation until iteration at $t_{a}=7000$, and Warm-Up until iteration at $t_{w}=7000$. For optimization, the weights for MS-SSIM and normal loss calculations are both set at $\lambda_{ms-ssim}=\lambda_{normal}=0.1$. When working with object-centered datasets that lack extensive backgrounds, we advise setting the scale learning rate $\eta_{s}=0$ to maximize geometric accuracy. Specifically for the DTU dataset, we set Prune Threshold $\epsilon=0.5$, loss weight $\lambda_{ms-ssim}=1$, Atom Scale $\mathcal{S}=P_{10}(\bm{d})$, ant total training iterations at 7k.
@@ Appendix E Additional Results
Figure 9: Radiance Field Comparison on the Mip-NeRF360 Dataset.
@@
Figure 10: Mesh Comparison on the DTU and NeRF Synthetic Datasets [Aanæs et al.(2016)Aanæs, Jensen, Vogiatzis, Tola, and Dahl, Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]